Time to swing. The period T (time in seconds for one complete cycle) of a simple pendulum is related to the length L (in feet) of the pendulum by the formula 8T^2 = pie^2L. If a chld is on a swing with a 10 foot chain, then how long does it take to complete one cycle of the swing?
Answer by fractalier(6550)

Okay...we plug 10 in for L and find T...
8(T^2) = (pi^2)L
T^2 = (pi^2)(10) / 8
T^2 = 12.3245 and
T = 3.51 sec